OpenStudy (anonymous):

one leg of a triangle is 7 ft longer than the other leg . The hypotenuse is is 13ft . find the length of the three sides of the triangle.

OpenStudy (mertsj):



OpenStudy (mertsj):

Use the Pythagorean Theorem

OpenStudy (maheshmeghwal9):

we know triplet (12,5,13).thus,12 &5 are Answers

OpenStudy (mertsj):

\[x^2+(x+7)^2=13^2\]

OpenStudy (maheshmeghwal9):

u get x=5 & x+7=12.

OpenStudy (maheshmeghwal9):

i am with mertsj!

OpenStudy (anonymous):

so 5, and 12

OpenStudy (anonymous):

or do i need to put 5, and x+7=12 for the answer?

OpenStudy (maheshmeghwal9):

u only write p=5,b=12,h=13. k!.:)

OpenStudy (anonymous):

thanks

OpenStudy (maheshmeghwal9):

ur welcome